Please help me slove this word mixture problem.
One canned juice drink is 15% orange juice, another is 5% orange juice. How many liters of each should be mixed together in order to get 10L that is 14% orange juice?

Answer by robertb(5830)

0.15x + 0.05(10 - x) = 0.14*10,
0.15x + 0.5 - 0.05x = 1.4,
0.10x = 0.9,
x = 9 liters, the amount of 15% orange juice that must be mixed with
10 - x = 1 liter of 5% orange juice.